INFRARED SPECTROSCOPY OF CH_{3}F-(ORTHO-H_{2})_{n} CLUSTERS IN SOLID PARAHYDROGEN

\underline{KAZUKI YOSHIOKA}, and DAVID T. ANDERSON, Depertment of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

Infrared spectroscopic studies of CH_{3}F isolated in solid parahydrogen (99.99% purity) revaled multiple peaks in the ν_{3} C-F stretch region. While the observed fine structure was first thought to be due to rotational motion of the CH_{3}F, further studies of samples with elevated orthohydrogen (o-H_{2}) concentrations clearly indicate the observed peaks are due to CH_{3}F-(o-H_{2})_{n} clustering. This has enabled infrared studies of CH_{3}F-(o-H_{2})_{n} (with n=1 to 12) clusters in solid parahydrogen. The observed shifts in the ν_{3} vibration with the number of o-H_{2} molecules clustering is similar to the shifts measured in (Ar)_{n}-HF clusters synthesized in helium nanodroplets^{a}. Experimental results in support of this interpretation will be presented including 2 ν_{3} and CD_{3}F studies.

aK. Nauta, R E. Miller, J. Chem. Phys. 115, 10138 (2001).